Nikolas borrowed $14,000 to purchase new equipment for his landscaping business. The bank has loaned him this money at an annual interest rate of 7.5%. How much will Nikolas owe the bank when he repays his loan in eighteen months using simple interest?
Answer #1

The simple interest formula is given by: A=P(1+r\times n)

where 'P' is the principal amount, 'A' the accumulated amount, 'r' is the rate of interest and 'n' is the duration (in year).

So,

A=P(1+r\times n)=14000(1+0.075\times \frac{18}{12})=15575

i.e. $15575

which is the amount he'd owe to the bank.